OpenStudy (anonymous):

Show all work in dividing the quantity y¨Csquared plus 5y minus 6 all over the quantity y¨Csquared minus 3y minus 10 by the quantity y¨Csquared plus 2y minus 3 all over the quantity y¨Csquared minus 4y minus 5 .Use complete sentences to explain how to simplify this expression. Can someone answer this using sentences??

OpenStudy (saifoo.khan):

\[\frac{y^2+5y-6}{y^2-3y-10} \times \frac{y^2+2y-3}{y^2-4y-5}\]Like this?

OpenStudy (anonymous):

You are dividing it. \[y^2+5y-6/y^2-3y-10 \div y^2+2y-3/y^2-4y-5\]

OpenStudy (saifoo.khan):

\[\frac{y^2+5y-6}{y^2-3y-10} \div \frac{y^2+2y-3}{y^2-4y-5}\]Now?

OpenStudy (anonymous):

Yes

OpenStudy (saifoo.khan):

Okay. So now first step will be the changing of division sign, which will flip the right fraction.\[\frac{y^2+5y-6}{y^2-3y-10} \times \frac{y^2-4y-5}{y^2+2y-3}\] Agree?

OpenStudy (saifoo.khan):

Now. we'll try to factor all the expressions.

OpenStudy (saifoo.khan):

\[\frac{(y-1)(y+6)}{(y-5)(y+2)} \times \frac{(y-5)(y+1)}{(y-1)(y+3)}\]

OpenStudy (saifoo.khan):

Can you solve now?

OpenStudy (anonymous):

I'm still a little bit lost.

OpenStudy (saifoo.khan):

Why? Where?

OpenStudy (anonymous):

I just dont know what to do. Im learning this online for home schooland just dont understand.

OpenStudy (anonymous):

Thank you!!

OpenStudy (saifoo.khan):

One you see it, see more examples from left on that same page. Then try your question. I'll check your working!